John’s Custom Computer Shop (JCCS) assembles computers for both individual and corporate customers. The company is organized into two divisions: Personal and Business. Once a computer is built, it is shipped to the customer. Billing for all customers is handled by the corporate Accounts Receivable Department. Accounts Receivable performs two major activities: billing and dispute resolution. Billing refers to preparing and sending the bills as well as processing the payments. Dispute resolution occurs when a customer refuses to pay, usually due to an error in billing.

The costs of the Accounts Receivable Department are allocated to the two divisions based on the number of bills prepared. Kyle, the manager of the Business division, has complained that the allocated costs from Accounts Receivable are beginning to make the business division look unprofitable and has asked you to recommend some changes to the allocation system. If he agrees with your recommendation, he will pass them on to the chief financial officer.

Data on costs and activities in the Accounts Receivable Department follow.

Personal Business Total
Number of bills prepared 600 350 950
Number of disputes 55 10 65

The Accounts Receivable Department incurred the following costs during the year.

Billing $ 52,250
Dispute resolution 35,750
Total $ 88,000

Required:

a. Under the current allocation system, what is the cost that will be allocated from Accounts Receivable to Personal? To Business?

b. Suppose the company implements an activity-based cost system for Accounts Receivable with two activities, billing and dispute resolution. What is the cost that will be allocated from Accounts Receivable to Personal? To Business? Use the number of bills prepared as the cost driver for billing costs and the number of disputes for dispute resolution costs.
Answer #1

a.

Current allocation system = Total costs of accounts receivable / Number of bills prepared

Current allocation system = $88,000 / 950 = $92.63158

Personal = 600*$92.63158 = $55,579

Business = 350*$92.63158 = $32,421

b.

Activity-based cost system:

Billing = $52,250 / 950 = $55

Dispute resolution = $35,750 / 65 = $550

Personal = $33,000(600*$55) + 30,250(55*$550) = $63,250

Business = $19,250(350*$55) + 5,500(10*$550) = $24,750